# Flow Annealed Importance Sampling Bootstrap meets Differentiable Particle Physics

### v1

@@ 1 Introduction
Context.
In the advent of the high-luminosity phase at the Large Hadron Collider (LHC), significant speed-ups in the simulation software are required to analyze the increasing amount of data [1, 2, 3]. The simulated data points are compared to measured data from particle collisions to understand the underlying fundamental physics processes in more detail. One important step in the LHC simulation chain is the generation of samples (“events”) based on matrix elements (MEs), that is, unnormalized distributions $p(x)$ over the outgoing 4-momenta $x$ that are derived from a theoretical model. MEs can be evaluated analytically; however, sampling from them is hard since they can be high-dimensional and multi-modal. Additionally, they are defined on limited support originating from mass and momentum constraints and can exhibit divergences. Standard sampling algorithms such as MadGraph [4], SHERPA [5], and PYTHIA [6, 7] simplify MEs based on physics knowledge and employ adaptive Monte Carlo methods to approximate the distribution. More recently, machine learning-based surrogate models like normalizing flows have gained momentum to improve the efficiency of the sampling process [8, 9, 10, 11, 12, 13, 14, 15, 16]. As training normalizing flows with the maximum-likelihood loss relies on the costly generation of a large training data set, alternative approaches show potential that evaluate the distribution of interest directly during training [8, 9, 10, 11, 13, 14, 15, 16]. These methods evaluate the target distribution with samples from the normalizing flow, which requires gradients to be backpropagated through this distribution to update the flow parameters. Therefore, this training mode usually requires differentiable MEs, which have only recently been proposed [17] and employed [15] for normalizing flow training. A similar method, called Flow Annealed importance sampling Bootstrap (FAB) [18], also relies on the evaluation of a differentiable target density and has been developed to obtain samples from Boltzmann distributions of molecules. FAB uses Annealed Importance Sampling (AIS) with Hamiltonian Monte Carlo (HMC) transition steps to improve the quality of the normalizing flow samples towards the distribution of interest. The resulting AIS samples and their weights are used to train the normalizing flow. Running HMC requires a differentiable target distribution, and we are the first to perform HMC-based updates on MEs.
@@ 2 Method
A normalizing flow [19, 20, 21] is a density estimator that consists of a series of learnable, invertible transformations that construct a differentiable bijection between a simple, chosen base distribution and an expressive flow distribution $q_{\theta}(x)$.
@@ 2.1 Training
If samples $x\sim p(x)$ are cheaply available, the fKLD serves as a loss function: $D_{\mathrm{KL}}(p\;\|\;q_{\theta})=\mathbb{E}_{p}\left[\log(p(x)/q_{\theta}(x))\right]$ which can be simplified to the negative log-likelihood loss $\mathcal{L}_{\mathrm{fKLD}}=-\mathbb{E}_{p}\left[\log q_{\theta}(x)\right]=-\frac{1}{N}\sum_{i=1}^{N}\log q_{\theta}(x_{i})$ over $N$ data points. It can be shown that this loss function results in a density $q_{\theta}$ with mass-covering properties [18]. The flow distribution $q_{\theta}(x)$ is evaluated with samples from $p(x)$ and the gradient computation is straightforward: $\nabla_{\theta}\mathcal{L}_{\mathrm{fKLD}}=-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log q_{\theta}(x_{i})$.
@@
Another way of quantifying the difference between two distributions is via the rKLD, $D_{\mathrm{KL}}(q_{\theta}\;\|\;p)=\mathbb{E}_{q_{\theta}}\left[\log(q_{\theta}(x)/p(x))\right]$. This loss function has mode-seeking properties, meaning that it is not guaranteed that the optimized distribution $q_{\theta}(x)$ covers all modes of $p(x)$ [18]. To obtain a gradient for this loss function $\nabla_{\theta}\mathcal{L}_{\mathrm{rKLD}}=\nabla_{\theta}\mathbb{E}_{q_{\theta}}\left[\log(q_{\theta}(x)/p(x))\right]$, the gradient has to be propagated through the evaluation of the target distribution $p(x)$, since the samples $x\sim q_{\theta}(x)$ used for the estimation of the expectation value depend on the parameters $\theta$ themselves. This is visualized in Fig. 5 for illustration. As a result, $p(x)$ has to be differentiable [^1] One can prevent differentiating through the target distribution by mapping the flow samples (with stopped gradient) back through the flow in the reverse direction and evaluating the loss function in latent space [13]. However, this approach requires two subsequent applications of the flow transformations and is more costly..
@@
[…]
where fKLD corresponds to $\alpha\rightarrow 1$ and rKLD to $\alpha\rightarrow 0$ [22]. In FAB [18], $D_{\alpha=2}$ is chosen as a loss function since it minimizes the variance of the importance weights $w=p(x)/q_{\theta}(x)$—a desirable property of a well-performing density estimator $q_{\theta}(x)$—and the resulting distribution $q_{\theta}(x)$ has mass covering properties with respect to $p(x)$. Since samples from the flow might poorly fit the target distribution at the beginning of training, AIS [23] is employed to pass the samples through a chain of intermediate distributions $q_{1},...,q_{M-1}$ with HMC as a transition operator. The intermediate distributions $q_{i}$ are chosen to interpolate between the flow distribution $q_{0}=q_{\theta}$ and the AIS target $q_{M}=p^{2}/q_{\theta}$. Importance weights $w_{\mathrm{AIS}}$ can be computed for the AIS samples $x_{\mathrm{AIS}}$ which allow evaluating the surrogate loss function $\mathcal{S}(\theta)=-\mathbb{E}_{\mathrm{AIS}}\left[\bar{w}_{\mathrm{AIS}}\log q_{\theta}(\bar{x}_{\mathrm{AIS}})\right]$ where the bar over variables indicates that the gradient is not propagated through the AIS sequence in the backward pass. To reduce the cost of evaluating the target distribution and its gradient for each intermediate distribution and each HMC step in between, the AIS samples can be stored in a replay buffer. Pairs $\{\bar{x}_{\mathrm{AIS}},\bar{w}_{\mathrm{AIS}}\}$ are sampled from the buffer based on the importance weights to perform multiple gradient updates per iteration. Fig. 5 shows a schematic illustration of FAB; for further details, see Midgley et al. [18].
@@ 2.2 Differentiable Matrix Elements
Both training with rKLD and FAB requires gradients of the target distribution. Recent developments in HEP provide us with differentiable implementations of complex amplitudes [17, 24, 15]:
$\bm{\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}}$.
For the $\Lambda_{c}^{+}$ decays [25, 26, 27], the ComPWA package [24] utilizes the SymPy engine [28] to formulate symbolic amplitude models whose compute graph can be transformed to JAX [29] for efficient gradient computation. We select the $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$ decay to serve as a two-dimensional example due to its complex resonant structure resulting from the presence of multiple decay chains. The amplitude can be visualized in a Dalitz plot [30], shown in Fig. 6.
$\bm{e^{+}e^{-}\rightarrow t\bar{t},t\rightarrow W^{+}b,\bar{t}\rightarrow W^{-}\bar{b}}$.
As a higher-dimensional example, we choose the eight-dimensional $t\bar{t}$ matrix element which can be generated with MadJAX [17] and provides a differentiable implementation of the scattering amplitudes of MadGraph [4]. A differentiable non-linear transformation is applied to map the irregular phase space boundaries to the unit hypercube based on RAMBO [31]. It is not possible to map the distribution to infinite support since divergences can occur at the phase space boundary, which would result in non-zero probability mass at infinity.
@@ 3 Experimental Setup
For each of the aforementioned MEs and training methods, we train three models with different random seeds and average their results to improve reliability (see App. C for details).
@@
We evaluate the performance of the normalizing flows based on three metrics: (1) the fKLD evaluated on a test data set, (2) the importance sampling efficiency calculated from the importance weights, and (3) the integral estimate of $p(x)$. The fKLD, introduced in Section 2, quantifies the difference between the learned distribution $q_{\theta}(x)$ and the true target distribution $p(x)$ by evaluating the normalizing flow with the samples $x\sim p(x)$. If both distributions match almost everywhere, the KL divergence is zero.
@@
$\epsilon=\frac{1}{N}\left[\sum_{i=1}^{N}w_{i}\right]^{2}\Big{/}\left[\sum_{i=1}^{N}w_{i}^{2}\right]\,.$ (2)
We elaborate on its derivation and similarities to the unweighting efficiency—which is commonly employed in HEP—in App. B. The integral estimate $\bar{I}$ of the target distribution $p(x)$ is defined as
[…]
For the 2D example, the sampling efficiency $\epsilon$ and the integral estimate $\bar{I}$ are calculated for $10^{4}$ flow samples, while $10^{6}$ flow samples are used for the 8D ME. We compare our results for $\bar{I}$ with the physics-agnostic grid-based integral estimation method called VEGAS+ [34, 35] which is described in App. C in more detail. To illustrate the model performance, we provide histograms of flow samples for $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$ as well as a corner plot for the eight-dimensional ME in App. D. The performance metrics are summarized in Tab. 1 for both MEs. To make rKLD and FAB comparable, we report results based on training runs with the same number of target evaluations.
Since the target distribution can be costly to evaluate, we additionally show the trend for the importance sampling efficiency $\epsilon$ as a function of the number of target evaluations. Since fKLD training relies on a pre-computed training data set on which the normalizing flow can potentially be trained for an arbitrary number of epochs, we do not include it in this consideration. For rKLD, the ME is calculated once for each batch of flow samples, while the target density and its gradient need to be evaluated for every HMC step between intermediate AIS distributions for FAB. To reduce the number of target evaluations, FAB allows multiple gradient updates per iteration with samples from the replay buffer. However, depending on the buffer size, more target evaluations might be required at the beginning of training to fill up the buffer.

### v2

@@ Abstract
  • Submitted for review: February 2025, Accepted for publication: May 2025
Keywords: Normalizing Flows, Particle Physics, Differentiable Programming, Annealed Importance Sampling, Generative Modeling
  This is the Accepted Manuscript version of an article accepted for publication in Machine Learning: Science and Technology. IOP Publishing Ltd is not responsible for any errors or omissions in this version of the manuscript or any version derived from it. This Accepted Manuscript is published under a CC BY licence. The Version of Record is available online at 10.1088/2632-2153/addbc1.
@@ 1 Introduction
In the advent of the high-luminosity phase at the Large Hadron Collider (LHC), significant speed-ups in the simulation software are required to analyze the increasing amount of data [1, 2, 3]. The simulated data are compared to measured data from particle collisions to understand the underlying fundamental physics processes in more detail. One important step in the LHC simulation chain is the generation of samples (“events”) based on matrix elements (MEs). MEs describe the dynamical information contained in particle interactions. The ME $\mathcal{M}=\Braket{p_{1},...,p_{n}}{\mathcal{M}}{p_{a},p_{b}}$ quantifies the transition from an initial state with momenta $p_{a}$ and $p_{b}$ to a final state with $n$ outgoing particles described by their momenta $p_{1},p_{2},...,p_{n}$. Using quantum field theory, MEs can be constructed by summing the contributions of all possible Feynman diagrams $\mathcal{M}=\mathcal{M}_{1}+\mathcal{M}_{2}+\mathcal{M}_{3}+...$ (also called “channels”). Individual MEs can be calculated using the Feynman rules [4], which can become computationally expensive to evaluate when higher-order terms are included. The dimensionality of the matrix element depends on the number of outgoing particles $n$, each having three spatial degrees of freedom. Taking mass and momentum constraints into account, the dimensionality of MEs amounts to $43n-$ 4. From a machine learning perspective, one can interpret MEs as unnormalized distributions $p(x)$ over the outgoing 4-momenta $\{p_{1},...,p_{n}\}$ which we will denote as $x$ in the following to simplify notation. Via the Feynman rules, MEs can be evaluated analytically; however, sampling from them is hard since they can be high-dimensional for large $n$ and multi-modal due to contributions from different channels. Additionally, they are defined on limited support originating from mass and momentum constraints and can exhibit divergences. It is possible to simplify the complicated multi-modal structure of MEs by decomposing the ME into its dominant channels, an approach referred to as multi-channeling [5].
Standard sampling algorithms such as MadGraph [6, 7], SHERPA [8], and PYTHIA [9, 10] rely on multi-channeling to reduce the complexity of the ME and employ adaptive Monte Carlo methods similar to VEGAS [11] to approximate the distribution of the individual channels. More recently, machine learning-based surrogate models like normalizing flows have shown to improve the efficiency of the sampling process compared to standard methods like VEGAS and multiple approaches have been proposed over the last years. In general, normalizing flows can be trained using divergence measures as loss functions that quantify the difference between the target and the flow distribution: Information about the target is either included by training with samples from the target or by evaluating the target distribution analytically. When training normalizing flows with the maximum-likelihood loss based on target samples [12], we have to rely on the costly generation of a large training data set. Therefore, approaches that evaluate the distribution of interest directly during training with samples from the flow, have been the focus in event generation: Most papers employ neural importance sampling (NIS) [13] which has been adopted to phase space sampling and combined with multi-channeling in the works of [14, 15, 16]. Building on these developments, the efficiency of this approach can be improved by introducing a replay buffer and a VEGAS based initialization, as well as adapting the coupling between dimensions in the normalizing flow [17, 18]. NIS without multi-channeling has been employed in the work of [19] where they extend the target to include a background density and develop a specific training scheme to boost the efficiency at the beginning of training. To make previous results accessible to non-specialist users, a dedicated library has been developed recently [20].
In this work, we take into account that gradients need to be backpropagated through the ME distribution to update the flow parameters when evaluating the target with samples from the flow. Therefore, this training mode usually requires differentiable MEs, which have only recently been proposed [21] and employed [22] for normalizing flow training. A similar method, called Flow Annealed importance sampling Bootstrap (FAB) [23], also relies on the evaluation of a differentiable target density and has been developed to obtain samples from Boltzmann distributions of molecules. FAB uses Annealed Importance Sampling (AIS) with Hamiltonian Monte Carlo (HMC) transition steps to improve the quality of the normalizing flow samples towards the distribution of interest. The resulting AIS samples and their weights are used to train the normalizing flow. Running HMC requires a differentiable target distribution, and we are the first to perform HMC-based updates on MEs. To compare the general performance of different training methods, we do not include domain-specific physics information via multi-channeling.
@@ 2 Method
A normalizing flow [24, 25, 26] is a density estimator that consists of a series of learnable, invertible transformations that construct a differentiable bijection between a simple, chosen base distribution and an expressive flow distribution $q_{\theta}(x)$. To generate samples $x\sim q_{\theta}(x)$, samples are obtained from the base distribution and passed through the subsequent transformations. The density $q_{\theta}(x)$ can be evaluated for a given data point $x$ by passing it through the inverse chain of transformations and evaluating the density of the base distribution.
@@ 2.1 Training
If samples $x\sim p(x)$ are cheaply available, the fKLD serves as a loss function:
$D_{\mathrm{KL}}(p\;\|\;q_{\theta})=\mathbb{E}_{x\sim p(x)}\left[\log\frac{p(x)}{q_{\theta}(x)}\right]~{}.$
It can be simplified to the negative log-likelihood loss
$\mathcal{L}_{\mathrm{fKLD}}=-\mathbb{E}_{x\sim p(x)}\left[\log q_{\theta}(x)\right]=-\frac{1}{N}\sum_{i=1}^{N}\log q_{\theta}(x_{i})$
over $N$ data points where we drop terms independent of $\theta$. It can be shown that this loss function results in a density $q_{\theta}$ with mass-covering properties [23]. The flow distribution $q_{\theta}(x)$ is evaluated with samples from $p(x)$ and the gradient computation is straightforward: $\nabla_{\theta}\mathcal{L}_{\mathrm{fKLD}}=-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log q_{\theta}(x_{i})~{}.$
@@
Another way of quantifying the difference between two distributions is via the rKLD,
$D_{\mathrm{KL}}(q_{\theta}\;\|\;p)=\mathbb{E}_{x\sim q_{\theta}(x)}\left[\log\frac{q_{\theta}(x)}{p(x)}\right]~{}.$
Compared to fKLD, we sample from the normalizing flow and evaluate the flow density $q_{\theta}(x)$ as well as the target distribution $p(x)$ with the obtained samples. As a result, $p(x)$ has to be analytically available which is the case for MEs. This loss function has mode-seeking properties, meaning that it is not guaranteed that the optimized distribution $q_{\theta}(x)$ covers all modes of $p(x)$ [23]. To obtain a gradient for this loss function
$\nabla_{\theta}\mathcal{L}_{\mathrm{rKLD}}=\nabla_{\theta}\mathbb{E}_{x\sim q_{\theta}(x)}\left[\log\frac{q_{\theta}(x)}{p(x)}\right]~{},$
the gradient has to be propagated through the evaluation of the target distribution $p(x)$, since the samples $x\sim q_{\theta}(x)$ used for the estimation of the expectation value depend on the parameters $\theta$ themselves. This is visualized in figure 4 for illustration. As a result, $p(x)$ has to be differentiable [^1] One can prevent differentiating through the target distribution by mapping the flow samples (with stopped gradient) back through the flow in the reverse direction and evaluating the loss function in latent space [17]. However, this approach requires two subsequent applications of the flow transformations and is more costly..
{subcaptionblock}
.5 Refer to caption {subcaptionblock}.5 Refer to caption
{subcaptionblock} Refer to caption
Figure 1: Training with fKLD
Figure 2: Training with rKLD
Figure 3: Training with FAB
Figure 4: Visualization of compared methods. The black arrows illustrate the forward pass, while the red arrows mark the backpropagation of the gradient to update the flow parameters $\theta$. For rKLD, the gradient propagation to the normalizing flow $q_{\theta}(x)$ (red) requires a differentiable target distribution $p(x)$ (blue). For FAB, the HMC updates compute the gradient of the distribution of interest, while the backpropagation of the gradient is stopped for the AIS samples $\bar{x}_{\mathrm{AIS}}$ and weights $\bar{w}_{\mathrm{AIS}}$. The dotted lines indicate that these steps do not occur when a gradient update is performed repeatedly with samples from the replay buffer.
@@
$D_{\alpha}(p\;\|\;q_{\theta})=-\frac{1}{\alpha(1-\alpha)}\int p(x)^{\alpha}~{}q_{\theta}(x)^{1-\alpha}~{}\mathrm{d}x~{},$ (1)
where fKLD corresponds to $\alpha\rightarrow 1$ and rKLD to $\alpha\rightarrow 0$ [27]. In FAB [23], $D_{\alpha=2}$ is chosen as a loss function since it minimizes the variance of the importance weights $w=p(x)/q_{\theta}(x)$—a desirable property of a well-performing density estimator $q_{\theta}(x)$—and the resulting distribution $q_{\theta}(x)$ has mass covering properties with respect to $p(x)$ [27]. Since samples from the flow might poorly fit the target distribution at the beginning of training, AIS [28] is employed to pass the samples through a chain of intermediate distributions $q_{1},...,q_{M-1}$ with HMC as a transition operator. The intermediate distributions $q_{i}$ are chosen to interpolate between the flow distribution $q_{0}=q_{\theta}$ and the AIS target $q_{M}=p^{2}/q_{\theta}$. For all HMC steps between two intermediate AIS distributions, the flow as well as the target distribution have to be evaluated and the gradient of both distributions is required to perform momentum updates. As a result, we require a differentiable target distribution for FAB with HMC. Importance weights $w_{\mathrm{AIS}}$ can be computed for the AIS samples $x_{\mathrm{AIS}}$ which allow evaluating the surrogate loss function
$\mathcal{S}(\theta)=-\mathbb{E}_{\mathrm{AIS}}\left[\bar{w}_{\mathrm{AIS}}\log q_{\theta}(\bar{x}_{\mathrm{AIS}})\right]~{},$
where the bar over variables indicates that the gradient is not propagated through the AIS sequence in the backward pass. To reduce the cost of evaluating the target distribution and its gradient for each intermediate distribution and each HMC step in between, the AIS samples can be stored in a replay buffer. Pairs $\{\bar{x}_{\mathrm{AIS}},\bar{w}_{\mathrm{AIS}}\}$ are sampled from the buffer based on the importance weights to perform multiple gradient updates per iteration. Figure 4 shows a schematic illustration of FAB; for further details, see [23].
@@ 2.2 Differentiable Matrix Elements
The training approaches explored and compared in this work are not limited to one area of event generation in particle physics. They only require a differentiable implementation of the matrix element, since training with rKLD and FAB requires gradients of the target distribution. Recent developments in HEP provide us with differentiable implementations of complex amplitudes [21, 29, 22]. We choose two examples of high interest for the particle physics community, one from flavor physics and one from collider physics:
$\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$.
In flavor physics, the resonances of baryons are studied to search for new states beyond the constituent-quark model [30, 31]. One recently analyzed example is the $\Lambda_{c}^{+}$ baryon and its decay into a proton, a kaon, and a pion [32, 33, 34]. This decay is particularly interesting since its resonance structure allows searches of CP symmetry violations [33] and new physics [35, 36]. When the $\Lambda_{c}^{+}$ baryon decays, it can do so via different decay channels, resulting in the $K^{-}\pi^{+}$, $pK^{-}$, and $p\pi^{+}$ systems. These decays are shown in figure 5 as Feynman-like diagrams.
Figure 5: The $\Lambda_{c}^{+}\rightarrow pK^{-}\pi^{+}$ decay (left-most diagram) consists of three decay channels characterized by their resonances $\Lambda_{0}$ (center-left), $\Delta^{++}$ (center-right), and $K^{*}$ (right), each visualized in Feynman-like diagrams.
Each of these systems decays via an intermediate particle that can have resonances which are shown in a Dalitz plot in figure 6. A Dalitz plot is a physics-specific visualization of 2D amplitudes where the axes are chosen such that the histograms can be interpreted as an unnormalized density [37]. The resonances appear in a Dalitz plot as vertical, horizontal, and diagonal structures. In figure 6, the vertical bands with two prominent lines correspond to the decay of $\Lambda\rightarrow pK^{-}$ resulting from the $\Lambda(1520)$ and the $\Lambda(1670)$ resonances. The significant horizontal band corresponds to the $K^{*}(892)$ meson based on $K^{*}\rightarrow K^{-}\pi^{+}$. Decays in $\Delta^{++}\rightarrow p\pi^{+}$ are visible on the diagonal of the Dalitz plot where the $\Delta^{++}(1232)$ resonance is notable. An example for interference effects among resonances of different decay channels is the horizontal $K^{*}(892)$ band that gets shifted when crossing the vertical line of the $\Lambda(1670)$ resonance. Destructive interference is visible at the upper corner of the Dalitz plot resulting from reciprocal influences of $\Lambda(1520)$ and higher-mass $K^{*}$ resonances [33]. The complex resonance structure of this decay is ideally suited for our study as it provides a low-dimensional, but challenging setting to compare the performance of different normalizing flow training methods. Additionally, the $\Lambda_{c}^{+}$ amplitude model is implemented in a differentiable way and the code is publicly available via the ComPWA package [29].
For the implementation of the $\Lambda_{c}^{+}$ decays, ComPWA utilizes the SymPy engine [38] to formulate symbolic amplitude models whose compute graph can be transformed to JAX [39] for efficient gradient computation. Since ComPWA was developed as a general purpose tool for analytically formulating amplitudes, researchers can implement new amplitude models. The specific shape of the Dalitz plot boundary originates from mass and momentum constraints. It would be suboptimal to train a normalizing flow directly on the Dalitz plot representation because the it could assign non-zero probability density outside the physically-plausible regions. To prevent this, we map the support of the ME to the unit interval in each dimension via a differentiable and invertible transformation which is explained in A. Similar transformations are applied in standard samplers [8, 40] and we employ rational-quadratic spline flows because they are explicitly designed for distributions with limited support. It is not possible to map the distribution to infinite support since kinematic divergences can occur at the phase space boundary [4], which would result in non-zero probability mass at infinity.
$e^{+}e^{-}\rightarrow t\bar{t},t\rightarrow W^{+}b,\bar{t}\rightarrow W^{-}\bar{b}$.
In collider physics, the $t\bar{t}$ reaction producing a $W$ boson and a bottom quark is repeatedly investigated due to the coupling of non-standard model processes to heavy elementary particles like the top quark [41, 42, 43, 44]. We employ $t\bar{t}$ production with unstable final states containing $W$ bosons from an electron positron collider as a challenging, higher dimensional example. We selected an $e^{+}e^{-}$ process since ML methods in event generation are relevant for future circular or linear electron-positron colliders [45, 46]. The eight-dimensional $t\bar{t}$ matrix element can be generated with MadJAX [21] which provides a differentiable implementation of the scattering amplitudes of MadGraph [6, 7]. With MadJAX, we can generate any leading-order amplitudes with spin $0,1/2$ or $1$ particles and the code can be extended to higher spin MEs. It is also possible to generate MEs for proton-proton collisions by including a differentiable implementation of the parton density [47, 22] which we leave for future work. To map the irrgeular phase space boundaries to the unit hypercube, a differentiable non-linear transformation based on the RAMBO algorithm is applied to the MadJAX ME [48]. The eight dimensions characterizing the final state of the selected ME can be interpreted as two rescaled, intermediate masses $\widetilde{M}_{1}$ and $\widetilde{M}_{2}$, as well as two angular variables $\cos\widetilde{\theta}_{k}$ and $\widetilde{\phi}_{k}$ with $k\in[1,2,3]$.
To summarize, both selected examples are based on libraries which allow users to generate differentiable implementations of arbitrary MEs. This means that the compared approaches of training normalizing flows can easily be applied to different MEs.
@@ 3 Experimental Setup
For each of the aforementioned MEs and training methods, we train three models with different random seeds and average their results to improve reliability. For details about the flow hyperparameters and train settings, we refer to B for details.
Baseline.
We compare our results with a physics-agnostic integral estimation method called VEGAS+ [11, 49]. This grid-based optimization method divides the support into a regular, rectangular grid and estimates the integral contribution of each subspace. With this information, the grid is iteratively updated to focus on regions with large contributions and the value for the integral estimate is calculated as a weighted average over multiple runs. Through a combination of stratified and importance sampling, VEGAS+ is fast and efficient and can account for correlations between dimensions. We describe the VEGAS+ specific hyperparameters and tuning procedure in detail in B. It is not possible to use MadGraph as a baseline for both reactions since it is tailored to collider physics and the complicated helicity structure of the $\Lambda_{c}^{+}$ decay is not implemented.
@@
We first provide qualitative performance comparisons for both MEs by showing histograms and a corner plot of samples obtained from VEGAS+ as well as the normalizing flows. For the $\Lambda_{c}^{+}$ ME, we compare the results to samples obtained by rejection sampling with a uniform proposal distribution. For the $t\bar{t}$ reaction, we include the samples obtained from MadGraph as a reference. Both sets of samples correspond to the training data sets employed for fKLD training. Secondly, we evaluate the quantitative performance of the normalizing flows based on three metrics: (1) the fKLD evaluated on a test data set, (2) the importance sampling efficiency calculated from the importance weights, and (3) the integral estimate of $p(x)$. The fKLD, introduced in Section 2, quantifies the difference between the learned distribution $q_{\theta}(x)$ and the true target distribution $p(x)$ by evaluating the normalizing flow with the samples $x\sim p(x)$. If both distributions match almost everywhere, the KL divergence is zero.
@@
$\epsilon=\frac{1}{N}\left[\sum_{i=1}^{N}w_{i}\right]^{2}\left[\sum_{i=1}^{N}w_{i}^{2}\right]^{-1}\,.$ (2)
We elaborate on its derivation and similarities to the unweighting efficiency—which is commonly employed in HEP—in C. The integral estimate $\bar{I}$ of the target distribution $p(x)$ is defined as
$\bar{I}=\int p(x)~{}\mathrm{d}x=\int q_{\theta}(x)\frac{p(x)}{q_{\theta}(x)}~{}\mathrm{d}x\approx\frac{1}{N}\sum_{i=1}^{N}w_{i}~{}.$ (3)
For the 2D example, the sampling efficiency $\epsilon$ and the integral estimate $\bar{I}$ are calculated for $10^{4}$ flow samples, while $10^{6}$ flow samples are used for the 8D ME. We compare our results for $\bar{I}$ with VEGAS+. The performance metrics are summarized in table 1 for both MEs. To make rKLD and FAB comparable, we report results based on training runs with the same number of target evaluations. Since the target distribution can be costly to evaluate, we additionally show the trend for the importance sampling efficiency $\epsilon$ as a function of the number of target evaluations. We do not include fKLD training in this consideration because it relies on a pre-computed training data set on which the normalizing flow can potentially be trained for an arbitrary number of epochs. For rKLD, the ME is calculated once for each batch of flow samples, while the target density and its gradient need to be evaluated for every HMC step between intermediate AIS distributions for FAB. To reduce the number of target evaluations, FAB allows multiple gradient updates per iteration with samples from the replay buffer. However, depending on the buffer size, more target evaluations might be required at the beginning of training to fill up the buffer.